OpenStudy (anonymous):

Write an equation of the line in slope-intercept form given the slope is -5 and the y-intercept is 7.

OpenStudy (anonymous):

just plug-in the values to form\[y=mx+b\]where m is the slope b is the y-intercept

OpenStudy (anonymous):

then simplify the equation

OpenStudy (anonymous):

soo is it Y=-5X+7

OpenStudy (anonymous):

yup you're correct...

OpenStudy (anonymous):

thx so much can u meh with more ques I have a test

OpenStudy (anonymous):

when is your test?

OpenStudy (anonymous):

i got 2 finish it asap :/ can u help me pls *-*

OpenStudy (anonymous):

i got a D tha first time i need 2 pull up my grade

OpenStudy (anonymous):

just put in your question...

OpenStudy (anonymous):

ok Write an equation of the line in slope-intercept form given the slope is 10 and the y-intercept is -3.

OpenStudy (anonymous):

the same procedure above...

OpenStudy (anonymous):

as long as the slope and y-intercept is given, just plug-in them in that formula, that is what we call Slope-Intercept form of a LINE\[y=mx+b\]

OpenStudy (anonymous):

sooo is it Y=10X-3

OpenStudy (anonymous):

yup you got it correct again!...\(\ddot\smile\)

OpenStudy (anonymous):

ok here i another it will take a min

OpenStudy (anonymous):

ok...i'll wait...

OpenStudy (anonymous):

Referring to the figure, write an equation of the line shown in the graph.

OpenStudy (anonymous):

you have two points already as P1(0,1) and P2(-1,0)... you may compute the slope of a line first using\[m=\frac{rise}{run}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

soo is it Y = -1X+5

OpenStudy (anonymous):

... then look for a point where x=0, its y is the y-intercept...

OpenStudy (anonymous):

your slope is -1... we can check...\[m=\frac{0-1}{-1-0}=\frac{-1}{-1}=+1\]

OpenStudy (anonymous):

Soo is it Y=1x+5

OpenStudy (anonymous):

one thing more about slope, a negative (-) slope is inclined to the left while a positive (+) slope inclined to the right...

OpenStudy (anonymous):

no still not... try to plug-in any of the given point P1... to give you the y-intercept... point P1 is actually the one you're looking for... it is a point where x=0, so it's y is the y-intercept....

OpenStudy (anonymous):

hmmm let meh think here

OpenStudy (anonymous):

our P1 is (0,1)... what is y here?

OpenStudy (anonymous):

remember that coordiantes of a point is in a form (x,y)...

OpenStudy (anonymous):

(-1,0)

OpenStudy (anonymous):

that is our point P2

OpenStudy (anonymous):

where x=-1 and y=0....

OpenStudy (anonymous):

in point P1 (0,1)... x=0 and its y=1 (this is the y-intercept)... Y-intercept of a line can be solve once x=0...

OpenStudy (anonymous):

therefore you have m=1 and b=1, what do you think is the line equation?

OpenStudy (anonymous):

im sorry Orion my internet is actin up

OpenStudy (anonymous):

no problem... just waiting here

OpenStudy (anonymous):

same procedure as before... you have slope m and y-intercept b...

OpenStudy (anonymous):

if I don't respond fast @Orion1213 then know that my internet is acting up

OpenStudy (anonymous):

yup i know... this one is much easier than the two you have just answered...

OpenStudy (anonymous):

soo is it Y=1x+1

OpenStudy (anonymous):

or you may simply write it as\[y=x+1\]if the coefficient is 1, we simply do not write it before its variable x, it is automatically 1... the same as the y on the left-hand side of the equation...

OpenStudy (anonymous):

still you got it correct...

OpenStudy (anonymous):

@Orion1213 im here u just cant c me n heres tha next ques Write a linear equation to model the situation. You borrow $70 from your brother. To repay the loan, you pay him $7 per week. Use w to denote the number of weeks and y as the balance on the loan.

OpenStudy (anonymous):

is it: y=7w

OpenStudy (anonymous):

is it hard orion

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

 from my drawing... our y-intercept is $70, that is at week w=0, to repay the loan, it will be pay at $7 a week... therefore we have the following trend of the line... w=0, y=70 w=1, y=70-7=63 w=2, y=63-7=56 ... and so on...

OpenStudy (anonymous):

hmmm interesting sooo is it y=70-7w

OpenStudy (anonymous):

@Orion1213

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

yes you're right... if w=10, you're balance would be y=0...

OpenStudy (anonymous):

ok next ques Write an equation of the line that passes through the point (3,0) and has the slope, m = -2.

OpenStudy (anonymous):

is it Y=-2X+6

OpenStudy (anonymous):

in this question we have a point and a slope, we can use the point-slope formula of\[(y-y_1)=m(x-x_1)\]

OpenStudy (anonymous):

so is my answer rite Y=-2X+6

OpenStudy (anonymous):

Plug-in the given values... let us check\[(y-0)=-2(x-3)\]\[y=-2x+6\]you're right!\(\ddot\smile\)

OpenStudy (anonymous):

you're learning it...

OpenStudy (anonymous):

ok so next 1 Write an equation of the line that passes through the point (1,2) and has the slope, m = 2.

OpenStudy (anonymous):

just the same from the previous one....

OpenStudy (anonymous):

sorry i have to leave now... i have a visitor to face for a while... your answer must be y=4x...

OpenStudy (anonymous):

ok thx 4 all ur help ill try 2 find some1 else 2 help thx again @Orion1213!!!!!!!!!! :)